OpenStudy (nuccioreggie):

Solve the system of equations and choose the correct answer from the list of options. d + e = 15 −d + e = −5 Label the ordered pair as (d, e). (0, 0) (10, −5) (5, 10) (10, 5)

OpenStudy (nuccioreggie):

@whovianchick

OpenStudy (anonymous):

(10,5)

OpenStudy (whovianchick):

lol im smart one sec

OpenStudy (nuccioreggie):

lol

OpenStudy (anonymous):

I don't think it is :/

OpenStudy (whovianchick):

d + e = 15 −d + e = −5 -d+e=-5 e=d-5 d+e=15 d+d-5=15 2d=20 d=10 d+e=15 10+e=15 e=5 (10,5)

OpenStudy (anonymous):

I was right :D

OpenStudy (whovianchick):

I subtracted 5 from 15 instead of adding it my bad :)

OpenStudy (nuccioreggie):

thank yall

OpenStudy (anonymous):

lol that's fine and your welcome :)